ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

18. Teaching and Popularization of Mathematics

1991 MS Classification: 49, 68, 90 Kuragina, Elena, Mari State University, Russia Methodical support for a programming-methodical complex of study methods of optimization In given work methodical support for a programming-methodical complex of study of methods of optimization of the decision of extreme problems [1], on the basis of a system ODiS is submitted. ODiS - dialogue system of optimization of including four known group of methods (nondifferentiable cost functions, feasible direction methods, methods of penalty and barrer functions) is submitted. By work with these complex a following sequence of actions is recommended: trained is acquainted with a theoretical part under the text of lectures, further the student decides testy problems, on a example of which work of methods is illustrated. The training program is further applied, which accompanies with process of the decision of a problem by the messages, illustrating each step of algorithm of chosen method. Problems are chosen so that during work of a system evidently to demonstrate work of various methods by choice of appropriate parameters of uniform algorithm. Then individual task from chosen of bank of problems is executed. Reference: Izhutkin V. S., Petropavlovskii M. V. ODiS - the Nonlinear Programming Optimization Dialog System, Abstracts of the 9th Syberian School-Seminar ”Optimization Methods and its Using”, (1992),53.